OpenStudy (anonymous):

Hello I need help please Factor14^{4}+ 35^{3} -7^{2}

OpenStudy (anonymous):

\[14^{4}+ 35^{3} -7^{2} \]

OpenStudy (anonymous):

wait its \[14x^{4}+ 35x^{3} -7x^{2}\]

OpenStudy (anonymous):

im sorry

OpenStudy (phi):

can you factor the numbers into their prime factors? also, remember x^4 is short for x*x*x*x the idea is to factor each term, and find things (letters or numbers) that are in each term that you can "pull out"

OpenStudy (anonymous):

so Phi I need to find factors for 14 and 35 and 7

OpenStudy (phi):

that is a good step

OpenStudy (anonymous):

14x^4+35x^3-7x^2 \[=7x ^{2}(2x ^{2}+5x-1)\]

OpenStudy (anonymous):

so Phi for 14 its 1& 14 , 2&7 35 is 1&35 , 5&7 and 7 is 1&7

OpenStudy (anonymous):

Surijithayer I need a detailed report that is how I understand things better :o)

OpenStudy (phi):

14 is 2*7 (though 1*14 are factors, they are not prime factors) here is a list of the first few primes 2 3 5 7 11 13 17 19 23 29 the idea (an important idea) is any number can be broken up into multiples of *only* primes. you get \[ 2 \cdot 7 \cdot x \cdot x \cdot x \cdot x + 5 \cdot 7 \cdot x \cdot x \cdot x - 7 \cdot x \cdot x \]

OpenStudy (phi):

now look for the same letters and numbers to "pull out" notice all terms have a 7 and 2 x's

OpenStudy (anonymous):

what is a prime number anyway?

OpenStudy (phi):

a prime number can only be divided by itself and 1. example: 3 = 1*3 (not other factors) not prime: 6 = 6*1 and 2*3 (2*3 is the prime factors version)

OpenStudy (phi):

most people would not bother expanding the x's because you can use the exponent to know how many x's are in each term.... all terms have at least 2 x's.... so pull out x^2

OpenStudy (anonymous):

Oh okay I get it (for both) and then what comes next(the equation)?

OpenStudy (phi):